# MOSAIC: Masked Objective with Selective Adaptation for In-domain Contrastive Learning

### v1

@@ 4.2 Ablation Studies
MOSAIC-Stage2 ($\alpha$ 0.1) 76.032
MOSAIC-Stage2 ($\alpha$ 0.2) 81.336
MOSAIC-Stage2 ($\alpha$ 0.3) 88.116
MOSAIC-Stage2 ($\alpha$ 0.4) 86.794
MOSAIC-Stage2 ($\alpha$ 0.5) 67.708
All-Token MLM ($\alpha$=0.3) 63.995
All-Token MLM ($\alpha$=0.1) 66.092
All-Token MLM ($\alpha$=0.005) 77.768
All-Token MLM ($\alpha$=0.001) 78.753
@@
No Joint Objective. Finally, we assess the impact of the joint MLM+contrastive objective by removing the second stage entirely and training solely with the contrastive objective after vocabulary expansion (“Contrastive only”). As shown in Table 2, omitting the MLM stage results in a performance drop from 88.1 to 84.4, indicating that joint training with domain-restricted MLM provides a meaningful boost over contrastive adaptation alone.
@@ 5 Experiment on a Low-Resource Domain
Experimental Setup. As noted in Section 1, the Islamic domain is a low-resource area, especially for English-language data. In-domain data suitable for training sentence embedding models is extremely scarce. To address this, we constructed an in-domain training set by extracting semantically related verse pairs from Tafseer Ibn Kathir and applying consistency filtering with the gte-base model, resulting in 7,587 high-quality pairs. Further details on the data construction process are provided in Appendix B.
@@
Although the Islamic domain in English is characterized by an extremely limited amount of available training data, it is notable for having a dedicated evaluation dataset—unlike many other low-resource domains. Recent efforts by Malhas and Elsayed (2020) have created a verified high-quality Qur’anic Reading Comprehension Dataset (QRCD), which includes questions frequently asked within the Islamic domain. The answers provided are exhaustive, meaning all Qur’anic verses directly responding to the questions have been thoroughly extracted and annotated. To increase the size of the evaluation set, we combine the training and development splits, resulting in a total of 169 queries for testing. Although QRCD is originally in Arabic, we employ verified English translations to enable evaluation in the English language. For retrieval collections, we use the Sahih International English translation. [^4] https://tanzil.net/trans/ We compare our final model with three strong general-domain embedding models using NDCG@10 as the evaluation metric. The implementation details follow those used for the biomedical model, with the following modifications: we add 3k domain-specific tokens to the vocabulary. Consistent with our biomedical ablation, the experiment with $\alpha$ parameters produced similar trends, with $\alpha=0.3$ yielding the most stable and highest-performing results (see Appendix 5). The weights for the MOSAIC-ID (Islamic Domain) and Islamic-embed-model (trained solely using a contrastive approach) are initialized with nomic-embed-text-v1unsup.
@@
Results. The MOSAIC-ID-Stage3 model achieves the highest NDCG@10 score (36.8; Table 3), surpassing both general-domain baselines and the contrastive-only variant (Islamic-embed-model). All models exhibit considerable variation in per-query scores (Figure 2), reflecting the challenging nature of the dataset, but our model’s upper quartile and mean are both higher. Notably, the lower whisker for the MOSAIC-ID-Stage3 does not reach the minimum value of 0, whereas the lower whiskers for the general-domain models extend to 0. This indicates that our model makes fewer completely incorrect predictions (i.e., queries with NDCG@10 = 0), while the comparison models sometimes fail to retrieve any relevant results for certain queries. The upper whiskers are similar across all models, suggesting comparable best-case performance, but the reduction in low and zero scores for our model contributes to its higher overall mean NDCG@10. This performance gap can be attributed to differences in pretraining data: while biomedical content constitutes a measurable minority of large-scale pretraining corpora (Xiao et al., 2023; Wang et al., 2022b; Li et al., 2023; Nussbaum et al., 2024), Islamic domain texts are almost absent (typically less than 0.01%). This negligible coverage leaves general-domain models ill-equipped to capture the linguistic and conceptual nuances of Islamic texts, making domain adaptation essential for low-resource areas.
@@ 6 Conclusion
We present MOSAIC, a novel approach for domain adaptation of sentence embedding models by jointly optimizing MLM and contrastive objectives. Unlike standard domain adaptation methods, which are typically applied at the language modeling stage or after task-specific training via data augmentation, our method leverages a model-driven approach for domain adaptation after contrastive training. We achieve robust gains in both high-resource (biomedical) and low-resource (Islamic) domains, surpassing general-domain baselines even with limited in-domain data.
@@ Appendix A Consistency-based Filtering Procedure
To further ensure data quality and minimize false positive pairs, we employed a semantic filtering procedure using the gte-base model. Specifically, we first sampled up to 1 million candidate query–document pairs from the initial dataset. Each query and document was independently encoded into dense vector representations using the gte-base sentence embedding model.
@@ Appendix B Curating Passages for Training the Islamic Domain Model
Appendix C TRECCOVID analysis
However, as shown in Table 1, all of our models lag on the TRECCOVID dataset (Voorhees et al., 2021). Inspection of the TRECCOVID queries reveals that, alongside core biomedical and clinical questions, a significant fraction focuses on social or policy aspects of the pandemic (approximately 20%). Such queries, addressing societal impacts or interventions like school reopening, may fall outside the primary scope of biomedical corpora used for model adaptation. This mismatch in domain coverage could partly explain the observed underperformance. Moreover, recent large-scale analyses of PubMed using embedding-based atlases have shown that COVID-19 literature forms a uniquely isolated cluster in embedding space, with strong internal topical fragmentation, further challenging biomedical models (Kobak et al., 2024).
Appendix D Additional Ablations
Model Score
MOSAIC-Stage2 (alpha 0.1) 76.032
MOSAIC-Stage2 (alpha 0.2) 81.336
MOSAIC-Stage2 (alpha 0.3) 88.116
MOSAIC-Stage2 (alpha 0.4) 86.794
MOSAIC-Stage2 (alpha 0.5) 67.708
All-Token MLM, $\alpha$=0.3 63.995
All-Token MLM, $\alpha$=0.1 66.092
All-Token MLM, $\alpha$=0.005 77.768
All-Token MLM, $\alpha$=0.001 78.753
MOSAIC-Stage2 (mlm_prob 0.3) 49.871
MOSAIC-Stage2 (as the 3d stage) 70.540
Contrastive only 84.428
@@ Appendix D Additional Ablations
Appendix E Training Hyperparameters
@@
Table 6: Hyperparameters for training and finetuning sentence embedding models.

### v2

@@ 4.2 Ablation Studies
MOSAIC-Stage2 $\alpha$=0.1 76.032
MOSAIC-Stage2 $\alpha$=0.2 81.336
MOSAIC-Stage2 $\alpha$=0.3 88.116
MOSAIC-Stage2 $\alpha$=0.4 86.794
MOSAIC-Stage2 $\alpha$=0.5 67.708
All-Token MLM, $\alpha$=0.3 63.995
All-Token MLM, $\alpha$=0.1 66.092
All-Token MLM, $\alpha$=0.005 77.768
All-Token MLM, $\alpha$=0.001 78.753
MOSAIC-Stage2 (mlm_prob 0.3) 49.871
MOSAIC-Stage2 (as the 3d stage) 70.540
@@
No Joint Objective. We assess the impact of the joint MLM+contrastive objective by removing the second stage entirely and training solely with the contrastive objective after vocabulary expansion (“Contrastive only”). As shown in Table 2, omitting the MLM stage results in a performance drop from 88.1 to 84.4, indicating that joint training with domain-restricted MLM provides a meaningful boost over contrastive adaptation alone.
@@ 5 Experiment on a Low-Resource Domain
Experimental Setup. As noted in Section 1, the Islamic domain is a low-resource area, especially for English-language data. In-domain data suitable for training text embedding models is extremely scarce. To address this, we constructed an in-domain training set by extracting semantically related verse pairs from Tafseer Ibn Kathir and applying consistency filtering with the gte-base model, resulting in 7,587 high-quality pairs. Further details on the data construction process are provided in Appendix B.
@@
Although the Islamic domain in English is characterized by an extremely limited amount of available training data, it is notable for having a dedicated evaluation dataset—unlike many other low-resource domains. Recent efforts by Malhas and Elsayed (2020) have created a verified high-quality Qur’anic Reading Comprehension Dataset (QRCD), which includes questions frequently asked within the Islamic domain. The answers provided are exhaustive, meaning all Qur’anic verses directly responding to the questions have been thoroughly extracted and annotated. To increase the size of the evaluation set, we combine the training and development splits, resulting in a total of 169 queries for testing. Although QRCD is originally in Arabic, we employ verified English translations to enable evaluation in the English language (Pavlova, 2025). For retrieval collections, we use the Sahih International English translation. [^5] https://tanzil.net/trans/ We compare our final model with three strong general-domain embedding models using NDCG@10 as the evaluation metric. The implementation details follow those used for the biomedical model, with the following modifications: we add 3k domain-specific tokens to the vocabulary. Consistent with our biomedical ablation, the experiment with $\alpha$ parameters produced similar trends, with $\alpha=0.3$ yielding the most stable and highest-performing results (see Appendix 4). The weights for the MOSAIC-ID (Islamic Domain) and Islamic-embed-model (trained solely using a contrastive approach) are initialized with nomic-embed-text-v1unsup.
@@
Results. The MOSAIC-ID-Stage3 model achieves the highest NDCG@10 score (36.8; Table 3), surpassing both general-domain baselines and the contrastive-only variant (Islamic-embed-model). All models exhibit considerable variation in per-query scores (Figure 2), reflecting the challenging nature of the dataset, but our model’s upper quartile and mean are both higher. Notably, the lower whisker for the MOSAIC-ID-Stage3 does not reach the minimum value of 0, whereas the lower whiskers for the general-domain models extend to 0. This indicates that our model makes fewer completely incorrect predictions (i.e., queries with NDCG@10 = 0), while the comparison models sometimes fail to retrieve any relevant results for certain queries. The upper whiskers are similar across all models, suggesting comparable best-case performance, but the reduction in low and zero scores for our model contributes to its higher overall mean NDCG@10. This performance gap can be attributed to differences in pretraining data: while biomedical content constitutes a measurable minority of large-scale pretraining corpora (Xiao et al., 2023; Wang et al., 2022b; Li et al., 2023; Nussbaum et al., 2024), Islamic domain texts are almost absent (typically less than 0.01%). This negligible coverage leaves general-domain models ill-equipped to capture the linguistic and conceptual nuances of Islamic texts, making domain adaptation essential for low-resource areas (Pavlova and Makhlouf, 2024). The qualitative analysis on QRCD is provided in the Appendix E.
@@ 6 Conclusion
We present MOSAIC, a novel approach for domain adaptation of text embedding models by jointly optimizing MLM and contrastive objectives. Unlike standard domain adaptation methods, which are typically applied at the language modeling stage or after task-specific training via data augmentation, our method leverages a model-driven approach for domain adaptation after contrastive training. We achieve robust gains in both high-resource (biomedical) and low-resource (Islamic) domains, surpassing general-domain baselines even with limited in-domain data.
@@ Appendix A Consistency-based Filtering Procedure
To further ensure data quality and minimize false positive pairs, we employed a semantic filtering procedure using the gte-base model. Specifically, we first sampled up to 1 million candidate query–document pairs from the initial dataset. Each query and document was independently encoded into dense vector representations using the gte-base text embedding model.
@@ Appendix B Curating Passages for Training the Islamic Domain Model
Appendix C Training Hyperparameters
@@
Table 5: Hyperparameters for training and finetuning text embedding models.
Appendix D Computational Efficiency
Method Relative Compute Cost Notes Baseline: Contrastive FT 1.0x Supervised dual-encoder finetuning MOSAIC – Stage 1 (Tokenizer extension) 0x No GPU cost MOSAIC – Stage 2 (Domain tokens only)  2.0x MLM applied only to new domain tokens MOSAIC – Stage 2 (All tokens)  2.5x Upper bound: MLM applies to entire vocabulary MOSAIC – Total  3.0x Stage 1 + Stage 2 (domain tokens) + CT finetuning Full Domain MLM (DAPT) 10–15x Typical costs reported in prior work
Table 6: Relative compute cost comparison between MOSAIC stages and common domain adaptation baselines.
As summarized in (Table 6), the computational footprint of MOSAIC remains modest. The second stage introduces an overhead of approximately 2× relative to standard contrastive finetuning, while the full three-stage pipeline remains substantially cheaper than conventional domain-adaptive pretraining (DAPT). Notably, the naïve all-tokens MLM variant incurs higher computational cost, which highlights the efficiency benefits of restricting masked language modeling to newly introduced domain tokens. Overall, these results show that MOSAIC can achieve meaningful performance gains in specialized domains without requiring prohibitive computational resources.
Appendix E Qualitative Analysis
For the qualitative analysis (Table 7), we present examples from the BIOSSES dataset. In the first two cases, the MOSAIC-Stage2 model produces similarity predictions that are close to the gold scores, indicating well-calibrated semantic judgments. In contrast, the All-Tokens MLM model consistently overshoots, assigning higher similarity scores than warranted. In the remaining examples, MOSAIC-Stage2 also overestimates similarity, but to a noticeably smaller extent than the All-Tokens MLM baseline. Given the high density of biomedical terminology in BIOSSES, the All-Tokens model appears biased toward inflated similarity estimates. This behavior suggests that, during the joint MLM+contrastive stage, the All-Tokens approach does not sufficiently emphasize domain-specific token distinctions, leading to weaker fine-grained semantic discrimination across biomedical topics.
[…]
Table 7: Qualitative analysis examples from the BIOSSES dataset comparing MOSAIC-Stage2 with an all-token MLM baseline. Predicted scores indicate semantic similarity.
For the qualitative analysis in the Islamic domain, we focus on the retrieval task from the QRCD benchmark (Malhas and Elsayed, 2020). Many queries in QRCD are genuinely challenging even for human readers, as they require recognizing metaphorical expressions and implicit semantic correspondences rather than relying on surface lexical overlap. To illustrate the effect of domain adaptation, we analyze two particularly challenging queries and compare the top three retrieved passages across models. We present one case in which the proposed MOSAIC-Stage3 model succeeds while baseline models fail, and a second case in which MOSAIC-Stage3 also fails, highlighting limitations that remain shared across all systems.
For Query 308 (With what did Allah compare this life to?), MOSAIC-Stage3 successfully retrieves two of the three gold passages within its top-3 results, whereas the baseline models—including those shown in the table—fail to retrieve any gold passage within their top-10 rankings (we report only the top-3 for brevity). This query is particularly challenging because, although phrased broadly, it requires identifying specific metaphorical analogies used in the Qur’an to describe the transient nature of worldly life. The baseline models tend to gravitate toward generally related thematic verses (e.g., monotheism, devotion, or revelation), rather than capturing the intended comparative semantic structure. In contrast, MOSAIC-Stage3 correctly prioritizes passages that employ the canonical imagery of rain-fed vegetation to convey the temporary and illusory character of worldly existence.
For Query 350 (What are the verses in which healing is mentioned?), none of the models retrieve any gold passage within their top-10 ranked results (again, only the top-3 are shown). All retrieved passages are thematically distant from the target concept of healing, indicating a shared difficulty in linking the query to the appropriate Qur’anic contexts. Instead, the models gravitate toward broadly devotional or theological verses with only superficial topical or lexical relevance. This failure case illustrates that, even with domain adaptation, QRCD contains queries that demand highly specific semantic grounding—including recognition of rare thematic cues and uncommon lexical expressions—which remains beyond the current capabilities of all evaluated models.
Model Top Retrieved Passages (abridged) Query 308: With what did Allah compare this life to? MOSAIC (Stage 3) (1) (doc 600) “The worldly life resembles the plants that blossom by the help of the water which God sends from the sky. After a short time all of them fade away and the winds scatter them…” [GOLD]; (2) (doc 428) “The example of the worldly life is like the water sent down from the sky which becomes mixed with the earth’s produce that people and cattle consume. When the land becomes fertile and pleasant, people think that they have control over it…” [GOLD]; (3) (doc 1250) “Those whose good deeds will weigh heavier will live a pleasant life…” Islamic-embed-model (1) (doc 1263) “Say, ‘He is the only God; there is none equal to Him.’ ”; (2) (doc 585) “If all humans and jinn gathered to bring the like of this Qur’an…”; (3) (doc 1264) “Say, ‘I seek protection from the Lord of the Dawn…’ ” nomic-embed-text-v1 (1) (doc 987) “Have you seen the one who has chosen his desires as his lord? … ‘The only life is this worldly life.’ ”; (2) (doc 94) “Have you heard about the one who argued with Abraham about his Lord…”; (3) (doc 908) “We have given all kinds of examples for mankind…” Query 350: What are the verses in which healing is mentioned? MOSAIC (Stage 3) (1) (doc 1237) “After every difficulty there is relief; when you are free, strive and be devoted to your Lord.”; (2) (doc 166) “God wants to guide you and forgive you; He wants to relieve your burden, mankind was created weak.”; (3) (doc 1172) “This chapter is a reminder; those who seek guidance will find mercy from God.” nomic-embed-text-v1 (1) (doc 801) “These are the verses of wisdom, a guidance and mercy for the righteous.”; (2) (doc 477) “These are the illustrious verses revealed in Arabic so you may understand.”; (3) (doc 938) “A revelation from the Merciful God; verses fully expounded, with glad news and warnings.” Islamic-embed (1) (doc 1265) “Say: I seek protection from the Lord of mankind, against the evil of temptations.”; (2) (doc 2) “The path of those whom You blessed, not those astray.”; (3) (doc 1) “You alone we worship, and from You alone we seek help; guide us to the right path.”
Table 8: Qualitative analysis examples from QRCD.